### v1

# Axis-Aligned Document Dewarping

Abstract
Document dewarping is crucial for many applications. However, existing learning–based methods primarily rely on supervised regression with annotated data without leveraging the inherent geometric properties in physical documents to the dewarping process. Our key insight is that a well-dewarped document is characterized by transforming distorted feature lines into axis‐aligned ones. This property aligns with the inherent axis‐aligned nature of the discrete grid geometry in planar documents. In the training phase, we propose an axis-aligned geometric constraint to enhance document dewarping. In the inference phase, we propose an axis alignment preprocessing strategy to reduce the dewarping difficulty. In the evaluation phase, we introduce a new metric, Axis-Aligned Distortion (AAD), that not only incorporates geometric meaning and aligns with human visual perception but also demonstrates greater robustness. As a result, our method achieves SOTA results on multiple existing benchmarks and achieves $\boldsymbol{\sim}$34.5improvements on the AAD metric.
1 Introduction
The digitization of documents dramatically facilitates our lives by converting information on physical documents into electronic data that can be used with digital devices. However, due to paper deformation, scanners must be fixed in position for precise scanning and acquisition, so this method has specific limitations. With the widespread use of mobile phones and cameras, photography has become a primary tool for digitization. Nonetheless, images taken by handheld cameras often suffer from geometric distortions and uneven lighting, negatively impacting the user’s reading experience and downstream tasks like optical character recognition (OCR).
In early research, many studies defined this problem as surface reconstruction from different imaging devices, including multi-view images [33, 19], stereo cameras [34], and structured light cameras [1, 26]. After that, some studies have leveraged document priors in single-view images to conduct extensive research, such as document boundaries [2], text lines [13, 14]. However, these are affected by the low-level features of the detected prior knowledge.
In the years of the flourishing development of deep learning, convolutional neural networks began to learn 2D deformation field mapping from deformed images to expected dewarping results [23, 5, 39, 7, 24, 17, 16, 8, 12, 41]. Data-driven methods perform supervised learning dewarping under the guidance of various supervision signals with synthetic annotations [23, 36, 17], such as depth [5], geometric semantics [25], control points [40, 36, 41], layout [17].
Although prior methods perform well on simple deformations, they often fail on document images with complex deformations. This is mainly because previous studies rely on strong supervision signals, such as control points, which are effective but have no geometric meaning, and text lines, which are meaningful but difficult to extract and lack generalization. These studies do not effectively use the inherent geometric information in the paper documents, which is investigated and exploited in this paper.
Figure 1: The top row is the warped local document image, and the bottom is the expected dewarping result. The main change feature is that the feature line aligns with the axis as much as possible. The dotted lines are added for better visibility.
In this paper, we perform document dewarping using a lightweight convolutional neural network that predicts a 2D unwarping grid from the distorted image. On this basis, to make the network better learn the nature of document dewarping, as shown in the dewarping example in Fig. 1, we expect the feature lines in the dewarped document to be more horizontally or vertically aligned. We leverage the inherent axis alignment property of the discrete grid geometry in planar documents, design a corresponding axis-aligned geometric constraint, and integrate it into the network training process. This dramatically improves the dewarping performance of the network. In the inference stage, to improve the dewarping performance for images containing target documents in different proportions and rotation angles, we propose an axis alignment preprocessing strategy. This strategy simplifies the dewarping process by leveraging the position properties of grid points and the axis alignment properties of planar document geometry. In the evaluation stage, we propose a new evaluation metric, AAD, to automatically measure the axis alignment distortion of the dewarped results, which have specific geometric meaning and align with human visual perception.
Our contributions are as follows.
  • We introduce an axis-aligned geometric constraint, significantly enhancing the performance of learning-based document dewarping.
  • We propose an axis alignment preprocessing strategy in the inference phase which boosts the dewarping performance robustly.
  • We propose a new evaluation metric, AAD, that not only incorporates geometric meaning and aligns with human visual perception but also demonstrates enhanced robustness.
2 Related Work
Current research on document dewarping can be broadly categorized into traditional optimization methods and deep learning-based methods. As mentioned in the Sec. 1, traditional approaches to document dewarping primarily involve image processing and 3D reconstruction, preceding the advent of deep learning.
The research on image processing is to model the most commonly used low-level features extracted from document images, including letters/words, boundaries, spacing, and text lines, and to correct perspective and page curl distortion through local line spacing and cell shape, which can be formulated as an energy minimization optimization problem [35, 30, 14, 31]. In the recent research, Jiang et al. [12] applied deep learning to extract boundaries and document lines, followed by geometric constraint optimization. However, it is not easy to detect and perceive low-level features in deformed document images, which limits the application of methods based on image processing.
In 3D reconstruction research, variations in shadows from a single image or across different imaging devices, along with other input cues, have been utilized to reconstruct 3D surfaces [32, 37, 43, 19, 9]. Some studies model these surfaces as special types, such as cylindrical surfaces [3] and developable surfaces [26]. Recently, Luo et al. [22] treat the document as a developable surface and employ a recently proposed isometric constraint for modeling [11]. However, these methods rely on detecting document feature lines, resulting in long optimization times and limited generality.
In the research of deep learning, the current research has generated a large amount of data using rendering techniques and dataset collection, such as Doc3D [5], Inv3D [10], Simulated paper [17], UVDoc [36]. These data-driven approaches can be categorized into forward mapping methods [23, 18, 12] and backward mapping methods [5, 6, 7, 8]. To further improve optimization, some methods incorporate additional supervision signals, such as text lines, masks, boundaries, layouts, angles, and curvatures to further improve optimization [42, 7, 4, 16, 17]. However, extracting these additional cues can introduce extra noise and lack generalizability.
An effective approach is to fully exploit the inherent geometric properties of the document, which are independent of annotated labels and generally more robust. As Xie et al. [39] exploited the continuity property of document deformation by incorporating local fairing constraints into the displacement prediction process. Furthermore, the recent GSO-Net [38], based on grid surface modeling, achieves accurate and robust optimization by learning the intrinsic geometric properties.
In this paper, we study and leverage the intrinsic geometric properties for document dewarping.
3 Method
We adopt the convolutional neural network architecture from UVDoc [36] as the foundation for our document dewarping network. This fully convolutional deep neural network simultaneously predicts the document’s 3D grid mesh and corresponding 2D unwarping grid in a dual-task framework, which is inspired by the work of Wang et al. [39, 40]
In the following sections, Sec. 3.1 details the axis-aligned geometric constraint introduced during training, Sec. 3.2 analyzes the loss functions employed during network training, and Sec. 3.3 describes the preprocessing strategy for axis alignment during inference.
3.1 Axis-Aligned Geometric Constraints
Figure 2: Axis-aligned geometric constraints were enforced during document dewarping network training. The top row shows the predicted 2D unwarping grid (obtained from Grid-Net in UVDoc [36]), and the bottom row displays its transformation into UV space based on the ground truth, facilitating the computation of the corresponding axis-aligned geometric constraints loss along horizontal and vertical directions.
Fig. 2 illustrates the pipeline for computing the axis-aligned geometric constraint loss during training. Our approach takes advantage of the inherent geometric properties of planar paper, where the corresponding discrete grid geometry is aligned along the horizontal and vertical axes (as shown by $Q^{GT}$ in Fig. 2). This property aligns with the horizontal and vertical orientations of the document’s feature lines in image space. So, we can assess the axis alignment distortion of these feature lines in the dewarped image by computing the corresponding deviation of the unwrapped grid in planar space. However, since the current predictions are in image space, directly calculating this error is challenging.
To address this issue, we first employ an interpolation function to map the predicted 2D unwarping grid from image space $P$ to UV space $Q$. In the UV space, the deviation from horizontal and vertical alignment is effectively quantified by computing the position variance along each row and column. By integrating these axis-aligned geometric constraints into the training process, we can guide the network in learning to dewarp distorted document feature lines toward the axis-aligned direction. The following are the details.
Given a distorted document image, a deep convolutional network predicts a 2D unwrapped grid points with sizes $h\times w\times 2$. Each point on the grid can be represented as:
$P=\{p_{i,j}\}_{i=1,j=1}^{h,w},\quad p_{i,j}=(x_{i,j},y_{i,j}).$ (1)
During training, the corresponding ground truth grid is,:
$P^{GT}=\{p^{GT}_{i,j}\}_{i=1,j=1}^{h,w},\quad p^{GT}_{i,j}=(x^{GT}_{i,j},y^{GT}_{i,j}).$ (2)
In UV space, these ground truth grid are a uniform grid:
$Q^{GT}=\{q^{GT}_{i,j}\}_{i=1,j=1}^{h,w},\quad q^{GT}_{i,j}=(u^{GT}_{i,j},v^{GT}_{i,j}).$ (3)
To quantify the distortion of the prediction results, we distort the correspondence points in $Q^{GT}$ according to the relative grid‐position relationships in image space:
$q=f(p)=\mathrm{Interpolate}\bigl{(}q^{GT};\,p^{GT},\,P\bigr{)}$ (4)
Thus, each predicted 2D grid point is transformed into the UV space as follows:
$Q=\{q_{i,j}\}_{i=1,j=1}^{h,w},\quad q_{i,j}=f(p_{i,j})=(u_{i,j},v_{i,j}).$ (5)
In UV space, the axis-aligned error of 2D grid $Q$ consists of two components. The horizontal error is quantified by the variance of the $v$ values along each row, and the vertical error by the variance of the $u$ values along each column, computed as follows:
$\mathcal{L}_{\text{hor}}=\sum\nolimits_{j=1}^{h}\mathrm{Var}\Big{(}\{v_{1,j},v_{2,j},\dots,v_{w,j}\}\Big{)},$ (6)
$\mathcal{L}_{\text{ver}}=\sum\nolimits_{i=1}^{w}\mathrm{Var}\Big{(}\{u_{i,1},u_{i,2},\dots,u_{i,h}\}\Big{)}.$ (7)
The overall axis-aligned geometric constraint loss is then defined as:
$\mathcal{L}_{\text{AL}}=\mathcal{L}_{\text{hor}}+\mathcal{L}_{\text{ver}}.$ (8)
Minimizing $\mathcal{L}_{\text{AL}}$ means the predicted 2D unwarping grid after being mapped into the UV space, as much as possible to align with the axis. Thus, the feature lines on distorted image dewarping results are axis-aligned as much as possible.
3.2 Loss Function
Let $I\in\mathbb{R}^{H\times W\times 3}$ be the input image. The network predicts 2D unwarping grid $G_{2D}\in\mathbb{R}^{h\times w\times 2}$ and corresponding 3D grid mesh $G_{3D}\in\mathbb{R}^{h\times w\times 3}$. The dewarped image $I_{d}$ is obtained via an interpolation and grid sampling using $I$ and $G_{2D}$. Our overall training loss is composed of the following components:
Grid 2D and 3D Losses. Following UVDoc [36], we supervise the network with the losses:
$\mathcal{L}_{2D}=\|G_{2D}-G_{2D}^{\text{gt}}\|_{1},\mathcal{L}_{3D}=\|G_{3D}-G_{3D}^{\text{gt}}\|_{1}$ (9)
where $G_{2D}^{\text{gt}}$ and $G_{3D}^{\text{gt}}$ denote the ground truth 2D grid and 3D coordinates, respectively.
Axis-Aligned Geometric Constraint Loss. We impose axis-aligned geometric constraints $\mathcal{L}_{AL}$ in the UV space, guiding the network to achieve effective axis alignment dewarping of feature lines, as depicted in Sec. 3.1.
SSIM Loss. Since geometric dewarping may introduce artifacts (e.g., shadows) in the dewarped images, a pixel-level MSE loss might over-constrain model optimization and lead to instability. Hence, we incorporate an SSIM (Structural Similarity Index Measure) loss to measure the similarity between the dewarped image and the ground truth, capturing both local structure and global consistency, which helps the model converge effectively:
$\mathcal{L}_{\text{SSIM}}=1-\mathrm{SSIM}(I,I_{d}).$ (10)
The overall loss is defined as:
$\mathcal{L}_{\text{all}}=\alpha\,\mathcal{L}_{2D}+\beta\,\mathcal{L}_{3D}+\gamma\,\mathcal{L}_{AL}+\lambda\,\mathcal{L}_{\text{SSIM}}.$ (11)
with hyperparameters set as $\alpha=\beta=1,\gamma=0.2,\lambda=0.05$, in our experiments.
3.3 Axis Alignment Preprocessing for Inference
During inference, previous methods employ additional segmentation and deformation models for preprocessing [24, 7] to improve dewarping robustness under diverse input conditions. Essentially, these methods distinguish between the background and the document region, enabling the model to focus on the document area and reducing dewarping difficulty. However, they do not directly address the elimination of geometric distortion.
Drawing on the insights from Fig. 1, axis alignment is central to the dewarping process. So, we propose an axis alignment preprocessing strategy, as illustrated in Fig. 3. For a given input image, we first compute the minimum‐area rotation rectangle using the positional information from the predicted 2D unwarping grid. Subsequently, we perform an axis alignment rotation and clip the target area to obtain a preprocessed result. We then reapply the inference process on this preprocessed image to yield a refined dewarping outcome. If necessary, this process can be iteratively repeated to achieve further improvement under varying conditions. As shown in Fig. 3, the fine results obtained after preprocessing outperform the coarse results, effectively reducing the influence of the target position on the overall correction.
Figure 3: Illustration of the dewarping inference process with axis alignment preprocessing. The red grid represents the predicted 2D unwarping grid, while the blue rectangle indicates the minimum-area rotated rectangle computed from it.
In contrast to previous preprocessing methods [24, 7], our approach eliminates the need for additional segmentation or deformation models. Instead, it directly leverages the predicted 2D unwarping grid to extract object region information, rendering the method more efficient and streamlined. A detailed analysis is provided in Sec. 4.5.
4 Experiments
The experimental section introduces the training data, test benchmarks, evaluation metrics, and implementation details, subsequently, we analyze the experimental results both quantitatively and qualitatively.
4.1 Datasets and Benchmarks
4.1.1 Training Datasets
We used two datasets, Doc3D [5] and UVDoc [36], for alternative optimization models following the setting of Verhoeven et al. [36].
Doc3D [5] is the largest distortion dewarping dataset, containing 100K distorted images. It was created using approximately 4,000 authentic document images and rendering software. The labels for each distorted document image include 3D coordinate maps, albedo maps, normals, depth maps, UV maps, and backward mapping maps.
UVDoc [36] aims to reduce the domain gap between the synthetic Doc3D dataset [5]—commonly used for training document unwarping models—and real document photographs. This dataset contains 20K pseudo-photorealistic document images and provides all the necessary information to train a coarse grid-based document unwarping neural network.
4.1.2 Benchmarks
We evaluated our method on the following three benchmarks:
DocUNet [23]: This benchmark comprises 130 distorted images captured in natural scenes using mobile devices. Following previous methods, 25 rich-text images were selected for OCR evaluation, and according to DocGeoNet [8], the 127th and 128th images (which did not match the labels) were rotated by 180.
DIR300 [8]: Captured with a moving camera, this benchmark contains 300 images of real distorted documents with complex backgrounds, varying degrees of distortion, and diverse lighting conditions. In line with previous studies, 90 rich-text images were selected for OCR evaluation.
UVDoc [36]: Generated via a pseudo-photorealistic rendering pipeline, this benchmark consists of 50 images, including matching pairs with and without illumination. Following the UVDoc protocol, 20 rich-text images were selected for OCR evaluation.
4.2 Evaluation Metrics
We evaluate our approach using image similarity metrics and optical character recognition (OCR) performance.
MS-SSIM, LD and AD. Following Ma et al. [23], we employ multi-scale structural similarity (MS-SSIM) and local distortion (LD) to assess image similarity. Additionally, we use the aligned distortion (AD) metric introduced in [24] to overcome the limitations of earlier measures.
CER and ED. Consistent with previous work, we measure OCR performance using Character Error Rate (CER) [27] and Edit Distance (ED) [15]. Our experiments utilize OCR recognition with Tesseract v5.4.0.20240606 and PyTesseract v0.3.13  [^1] https://pypi.org/project/pytesseract/.
Subsequently, we introduce the newly proposed metric.
4.2.1 AAD Metric
As illustrated in Fig. 1, our visual evaluation of document dewarped primarily relies on the axis alignment error of the document’s feature lines, which carries clear geometric significance. However, existing metrics fail to capture this characteristic effectively. We introduce the AAD (Axis-Aligned Distortion) metric, which quantifies distortion by measuring the axial alignment error of the dewarped results. The metric meaning is the same as in 3.1, the key difference is that this metric is computed during the evaluation phase using given image pairs rather than 2D grid pairs from the training phase.
Let $y$ denote the ground truth image. The optical flow field from $y$ to its dewarped version is computed using the SIFT-flow algorithm [20], yielding flow components $v_{x}$ and $v_{y}$ along the horizontal and vertical directions, respectively. The directional gradients of $y$ are then extracted using the Sobel operator [29] along each axis and normalized:
$\tilde{g}_{i}=\frac{\left|\text{Sobel}_{i}(y)\right|}{\max\left(\left|\text{Sobel}_{i}(y)\right|\right)},\quad i\in\{x,y\}.$ (12)
For each row $i$ and column $j$, we compute the gradient-weighted means and the corresponding deviations as follows:
$m_{i}=\frac{\sum_{j}v_{y}(i,j)\,\tilde{g}_{y}(i,j)}{\sum_{j}\tilde{g}_{y}(i,j)+\epsilon},d^{\text{row}}_{i,j}=\tilde{g}_{y}(i,j)\cdot\left|v_{y}(i,j)-m_{i}\right|,$ (13)
$n_{j}=\frac{\sum_{i}v_{x}(i,j)\,\tilde{g}_{x}(i,j)}{\sum_{i}\tilde{g}_{x}(i,j)+\epsilon},d^{\text{col}}_{i,j}=\tilde{g}_{x}(i,j)\cdot\left|v_{x}(i,j)-n_{j}\right|.$
Here, $m_{i}$ and $n_{j}$ denote the gradient-weighted mean of the vertical flow in the $i$th row and the horizontal flow in the $j$th column, respectively, while $d^{\text{row}}_{i,j}$ and $d^{\text{col}}_{i,j}$ measure the corresponding deviations in pixel ($i\,j$).
Define the pixel-wise overall deviation as the Euclidean combination of the row and column deviations, and let the AAD be their average over all $N$ pixels:
$d_{i,j}=\sqrt{\left[d^{\text{row}}_{i,j}\right]^{2}+\left[d^{\text{col}}_{i,j}\right]^{2}},\text{AAD}=\frac{1}{N}\sum_{i,j}d_{i,j}.$ (14)
The AAD metric uses the normalized image direction gradient weight to calculate the axial alignment error of the feature line in the row and column, lower values indicate a better dewarping effect.
Fig. 4 shows an example visualization of the AAD metric. For two different dewarped results, the alignment of table feature lines and the horizontal alignment of text lines serve as the primary evaluation criteria. By overlaying the horizontal and vertical components of the AAD heatmap onto the dewarped results, the corresponding axis alignment errors are displayed in detail. The heatmap colors correlate well with the calculated values, offering a perceptual measure of axis alignment.
Figure 4: Example of visualizing AAD metrics. The first row is the ground truth image and its gradient heatmap in the horizontal ($G_{y}$) and vertical ($G_{x}$) directions; the second and third rows correspond to the two different dewarped results and their overlay heatmaps for AAD and its horizontal (AAD_H) and vertical (AAD_V) components, it can according to the $\updownarrow$ and $\leftrightarrow$ arrows to see the distortion error in the alignment direction of the corresponding axes.
To validate the robustness of the proposed AAD metric, we applied geometric and illumination disturbances to the dewarped images shown in Fig. 5, generating 100 warped images for each setting. The known disturbance serves as the optical flow information for calculating the ground truth metrics, while the optical flow information obtained via the SIFT-flow[20] method is used to compute the reference metrics. The linear correlation between the reference value and the actual value of each metric can be a criterion for their robustness. As shown in Fig. 6(a), the AAD metric demonstrated the highest linear correlation ($R^{\text{2}}$) across all sets, indicating a strong linear relationship with the ground truth even under minimal geometric disturbances(i.e., in regions with lower metric values, which are critical for comparing dewarped results that closely approximate the true values), In contrast, the AD metric exhibited a nonlinear response in these regions. In Fig. 6(b), we evaluated the normalized standard deviation under three different illumination conditions for each image in Fig. 5. The results revealed that the standard deviation for AAD (0.0230) was smaller than that for AD (0.0252), further confirming its enhanced robustness.
Figure 5: Geometric and illumination disturbances applied to the dewarped image. On the left is the distorted corresponding to the dewarped and ground truth images. On the right, geometric and illumination disturbances are added to the dewarped image, and the intensity increases along with the corresponding arrow direction. Set1 includes only geometric disturbances, Set2 adds color interference to Set1, and Set3 adds a random shadow to Set2.
Figure 6: Robustness analysis of each metric. (a) The x-axis represents the ground truth metric computed from the ground truth optical flow, while the y-axis represents the reference metric derived from the SIFT-flow method; ideally, they should exhibit a high linear correlation. (b) The normalized standard deviation of the metrics under three different environmental conditions, lower values indicate greater robustness to environmental disturbances.
4.3 Implementation Details
Our model is implemented using the PyTorch framework [28]. The dewarping model is trained on two datasets: Doc3D [5] (88k images) and UVDoc [36] (20k images). The input images are resized to 712$\times$488 pixels, while the output 2D grid is set to 45$\times$31. We optimize the model using the AdamW optimizer [21] with a batch size of 72. The initial learning rate is $1\times 10^{-5}$ and is linearly decayed to 0 during training. The network is trained for 30 epochs on two NVIDIA 4090 GPUs, which takes approximately 15 hours. During inference (see Sec. 3.3), a single round of preprocessing is applied for the DocUNet and UVDoc benchmarks, whereas for the DIR300 benchmark—with its smaller target objects—preprocessing is performed twice.
4.4 Experimental Results
We compare our method against existing approaches both quantitatively and qualitatively on the aforementioned benchmarks.
Table 1: Quantitative evaluation on the DocUNet benchmark [23]. We compare our method with existing methods in terms of Multi-Scale Structural Similarity (MS), Local Distortion (LD), Aligned Distortion (AD), Axis-Aligned Distortion (AAD), Character Error Rate (CER) and Edit Distance (ED), “↑” indicates the higher the better and “↓” means the opposite, Bold font denotes the best result, and underline indicates second-best. The “Improvements” row shows the relative gain of our metric over the previous best.
Method MS↑ LD↓ AD↓ AAD↓ ED↓ CER↓
DewarpNet[5] 0.474 8.362 0.398 0.164 824.5 0.225
DDCP[40] 0.472 8.982 0.428 0.159 1124.7 0.280
DocTr[7] 0.509 7.773 0.369 0.151 708.6 0.185
PaperEdge[24] 0.473 7.967 0.368 0.121 731.4 0.183
DocGeoNet[8] 0.504 7.717 0.381 0.158 695.0 0.187
FAM[16] 0.497 8.416 0.377 0.151 688.3 0.176
Layout[17] 0.525 6.706 0.300 0.121 689.8 0.180
UVDoc[36] 0.545 6.827 0.316 0.125 754.2 0.193
Ours 0.543 6.249 0.278 0.099 603.1 0.150
Improvements -0.4% 6.8% 7.3% 18.2% 12.4% 14.8%
[…]
Figure 7: Qualitative evaluation on the DocUNet benchmark. The figure shows the dewarped results along with the corresponding heatmap overlays of the AAD component metric. The horizontal ($\leftrightarrow$) and vertical ($\updownarrow$) arrows indicate the directions for visualizing the axis-aligned error of the dewarped feature lines.
In Tab. 1, we compare our method with previous approaches on the DocUNet benchmark. Our approach demonstrates significant improvements across most metrics and achieves performance comparable to UVDoc [36] on the MS metric. Notably, our method delivers an 18.2% improvement on the proposed AAD metric. Furthermore, we performed a qualitative analysis of global images (see Fig. 7) by selecting several test cases rich in horizontal and vertical features for dewarping comparison. To highlight the distortion areas more clearly, we overlay the dewarped results from different methods with heatmaps generated from the horizontal and vertical components of the proposed AAD metric (an example is provided in Fig. 4). For our method, the resulting heatmaps appear fainter on documents with feature lines in varying orientations, indicating a reduced axis alignment error. As expected from the dewarping results in Fig. 1, the feature lines exhibit improved axis alignment in the dewarped images.
Table 2: Quantitative evaluation on the DIR300 benchmark[8]. The evaluation protocol is identical to that described in Tab. 1.
Method MS↑ LD↓ AD↓ AAD↓ ED↓ CER↓
DewarpNet[5] 0.492 13.944 0.332 0.147 1076.8 0.336
DDCP[40] 0.552 11.02 0.357 0.132 739.3 0.257
DocTr[7] 0.616 7.189 0.255 0.107 698.4 0.211
PaperEdge[24] 0.584 7.968 0.255 0.091 447.4 0.177
DocGeoNet[8] 0.638 6.403 0.241 0.106 672.5 0.206
FAM[16] 0.607 7.680 0.243 0.108 639.4 0.198
Layout[17] 0.652 5.702 0.195 0.087 495.4 0.173
UVDoc[36] 0.621 7.730 0.219 0.101 614.0 0.237
Ours 0.702 4.261 0.131 0.057 405.8 0.132
Improvements 7.7% 25.3% 32.8% 34.5% 9.3% 23.7%
[…]
Figure 8: Qualitative evaluations on the DIR300 benchmark. We compare our method with existing methods as described in Fig. 7.
In Tab. 2, we evaluate our method on the DIR300 benchmark, which comprises 300 images—the largest among the three benchmarks. Our approach achieves notable improvements across all metrics, with the AAD metric improvement to 34.5%. Furthermore, the dewarped results significantly boost OCR performance, with the CER reaching 23.7%. These improvements are qualitatively demonstrated in Fig. 8. Our method substantially reduces the distortion of document feature lines, resulting in fainter heatmap colors and yielding more complete and straighter contours.
Table 3: Quantitative evaluation on the UVDoc benchmark [36]. The evaluation protocol is identical to that described in Tab. 1.
Method MS↑ LD↓ AD↓ AAD↓ ED↓ CER↓
DewarpNet[5] 0.589 5.666 0.193 0.085 297.8 0.117
DDCP[40] 0.585 11.582 0.290 0.109 251.8 0.094
DocTr[7] 0.697 8.146 0.161 0.075 110.8 0.034
DocGeoNet[8] 0.706 8.350 0.168 0.072 123.1 0.047
RDGR[12] 0.610 6.423 0.281 0.096 122.45 0.051
UVDoc[36] 0.785 6.810 0.119 0.052 122.7 0.046
Ours 0.800 5.834 0.098 0.040 104.3 0.036
Improvements 1.9% -3.0% 17.6% 23.1% 5.9% -5.9%
[…]
Figure 9: Qualitative comparisons on the UVDoc benchmark. We compare our method with existing methods as described in Fig. 7.
In Tab. 3, we present the quantitative results on the UVDoc benchmark. Our method outperforms competing approaches across all metrics except for LD and CER, where our results are very close to the best values. Notably, the AAD metric improves by 23.1%. In the qualitative comparison shown in Fig. 9, dewarped results in both horizontal and vertical directions are displayed. These comparisons reveal that our method achieves excellent axis-aligned dewarping performance even in regions without textual features, highlighting the superiority of our approach.
4.5 Ablation Study
We conducted ablation experiments to evaluate the impact of two key components: the axis-aligned geometric constraint used during training (Sec. 3.1) and the axis alignment preprocessing applied during inference(Sec. 3.3).
As summarized in Tab. 4, both components independently enhance performance across three benchmark datasets. Specifically, for the DocUNet and UVDoc benchmarks—where a large proportion of targets are present—the axis-aligned geometric constraint yields greater performance improvements than the axis alignment preprocessing. Conversely, in the DIR300 benchmark, which contains a smaller proportion of targets, the benefits of the preprocessing strategy are more pronounced. Importantly, the combination of both modules produces the best overall performance, demonstrating their complementary advantages in different scenarios.
Table 4: Ablation study on various benchmarks. We evaluate the impact of the axis-aligned geometric constraint loss (AL) and the axis alignment preprocessing strategy (AP) module. The symbols ✗ and ✓ indicate that the corresponding module is disabled or enabled, respectively. Bold font denotes the best performance, while underline values indicate the second-best scores.
[…]
DocUNet[23] 0.538 6.897 0.322 0.124 638.1 0.167
0.549 6.566 0.291 0.105 718.0 0.183
0.536 6.767 0.308 0.120 628.5 0.165
0.543 6.249 0.278 0.099 603.1 0.150
DIR300[8] 0.615 7.691 0.215 0.100 633.0 0.243
0.629 7.880 0.176 0.076 498.9 0.206
0.694 4.303 0.152 0.069 541.6 0.153
0.702 4.261 0.131 0.057 405.8 0.132
UVDoc[36] 0.794 6.356 0.118 0.052 120.4 0.041
0.785 5.903 0.116 0.044 105.2 0.035
0.790 6.024 0.111 0.049 118.2 0.043
0.800 5.834 0.098 0.040 104.3 0.036
Fig. 10 provides several examples from our ablation experiments. For targets with a smaller proportion (middle row), the preprocessing strategy improves target integrity, while the addition of the axis-aligned geometric constraint further enhances the dewarping performance. For targets with a larger proportion (first and last rows), the performance gains from the axis-aligned geometric constraint are more prominent. Overall, the optimal dewarped results are achieved by applying both components simultaneously, which aligns with the improvements observed in the qualitative metrics in Tab. 4.
[…]
Figure 10: Dewarped results under different ablation settings. From left to right, each row displays the distorted input image followed by the dewarped outputs generated by different ablation settings. The heatmap value is the horizontal component of the AAD metric.
5 Conclusion
This paper proposes a document dewarping method that exploits the inherent axis-aligned properties of planar documents’ discrete grid geometry. By incorporating an axis-aligned geometric constraint during network training and applying an axis alignment preprocessing step during inference, our method outperforms state-of-the-art methods across multiple benchmarks. Experimental results demonstrate that these components are both independently effective and complementary. We also introduce a novel metric, AAD, which offers a clear geometric interpretation, aligns well with human visual perception, and exhibits enhanced robustness. Unlike previous studies that primarily focus on network architecture or dataset contributions, our approach leverages the intrinsic geometry of documents to improve dewarping performance. Future work will further explore these geometric properties and extend their application to other image dewarping tasks.